OpenStudy (anonymous):

Ok I'm stuck 100x^2-49y^2 Is this prime ??

OpenStudy (ghazi):

100x^2=(10x)^2

OpenStudy (anonymous):

woooooow cant u people help me im stuck with my quest for over 10 min

OpenStudy (zzr0ck3r):

what is the exact question? Are you asking if it is in its prime factorization?

OpenStudy (zzr0ck3r):

And calm down...I would kill for a question that took only ten minutes:)

OpenStudy (anonymous):

it says factor the binomial completely, but then it has an option to click all ready prime. and I am stuck on the x and y portion.

OpenStudy (phi):

do you know that (a^2 - b^2) = (a+b)(a-b) ? are you terms perfect squares?

OpenStudy (anonymous):

yes??? because of the 100x^2 and the 49y^2??

OpenStudy (phi):

so expect it to factor as (10x-7y)(10x + 7y)

OpenStudy (anonymous):

oh ok shoot!! when other people do these problems it seems easy,,